### v1

# Algorithm Generation via Creative Ideation

Abstract
Designing system algorithms remains challenging, where the discontinuous nature of the solution space often forces system engineers to rely on generic heuristics at the expense of performance. We study whether LLMs can practically drive algorithm generation, and find that they are biased towards well-known generic designs, rather than making the creative leaps needed to navigate the discontinuous solution space. To address this limitation, we introduce MetaMuse, a framework for creative ideation built on three self-reflection principles: (1) quantifying solution diversity and usefulness in measurable performance space, rather than abstract idea space, (2) steering ideation through external stimuli, rather than internal randomness, and (3) constructing executable solutions using waypoint reasoning, rather than free-form chain-of-thought. Extensive evaluation shows that MetaMuse can generate high-performing solutions for two critical problems at a global cloud provider: cache replacement (reducing cache misses by up to 35.76%) and online bin packing (reducing bin usage by up to 30.93%).
1 Introduction
Designing system algorithms continues to be a central challenge in computing systems. Traditionally, the development of such algorithms has been a manual and labor-intensive process. Our experience at a global cloud provider indicates that even seemingly simple algorithms used in production — such as cache replacement for data storage or bin packing for job scheduling — can require tens of thousands of engineering hours to design. As a result, practitioners often resort to generic heuristics from the literature, e.g., least-recently used (LRU) for cache replacement and first-fit for bin packing, which frequently result in suboptimal performance.
This paper asks whether large language models (LLMs) can practically drive algorithm generation, with an emphasis on principles to transform this task into a systematic process. The core challenge in system algorithm design arises from the nature of its solution space: it is an inherently discontinuous space, where even a small change in algorithm design (e.g., data structure or control flow) can lead to sharp and non-linear changes in performance. Although it is sometimes possible to estimate the upper-bound performance, searching for practical solutions that approach this bound remains non-trivial. Furthermore, the discontinuous solution space does not provide sufficiently predictable patterns or a smooth landscape to guide the search.
Due to this discontinuity, we approach the algorithm generation task from a different angle, and frame it as a sampling process in the solution space. LLM attempts to generate distinct solutions at each step. This generative process represents a sequence of leaps in discontinuous solution space (Bubeck et al., 2023), which we formulate as creative ideation for LLMs. In fact, the systems community has long hypothesized algorithm design as a discovery process of ideas (Kant, 1985).
To study the algorithm generation task, we focus on high-impact problems at a global cloud provider: cache replacement and online bin packing. Our initial attempts of repeatedly sampling GPT-4o, Llama3.3-70B, and DeepSeek-V3 show that LLMs are fundamentally hindered by availability bias (Tversky & Kahneman, 1973) — LLMs are trained to output the most likely sequence of words, according to training datasets. As a result, solutions tend to cluster around well-known heuristics in the literature, e.g., least-recently (LRU) and least-frequently used (LFU) for caching. Furthermore, we find that this bias cannot be practically addressed through LLM hyperparameters like temperature (Ackley et al., 1985).
The key to creative ideation is exploiting knowledge that LLMs assume to be probabilistically irrelevant to the given problem. What is missing is a self-reflection process, which thinks how to generate subsequent solutions by inspecting what solutions have been generated so far. In realizing such a self-reflection framework, MetaMuse, we see three model-agnostic principles surfacing to best guide this self-reflection. First, evaluating diversity of generated solutions should be grounded in the measurable feedback space (e.g., simulation performance of system algorithms), rather than the abstract idea space (Sankar & Sen, 2024). Second, steering the ideation is achieved through external stimuli (e.g., keywords), rather than internal randomness (Honghua Chen and Nai Ding, 2023). Third, developing executable solutions from external stimuli takes structured checkpoint-based steps, or waypoint reasoning, rather than free-form chain-of-thought (Wei et al., 2023; Mehrotra et al., 2024).
Our work makes the following contributions:
1.
Empirical analysis of LLMs’ fundamental limitations in the algorithm generation task (§2) — shedding light on the impracticality of relying on LLMs’ internal randomness for ideation, based on two critical problems at a global cloud provider: cache replacement and online bin packing.
2.
Principled framework for creative ideation, MetaMuse (§3) — systematically combining three self-reflection principles, to guide the discontinuous leaps necessary to overcome LLMs’ bias.
3.
Practical evaluation of the algorithm generation task (§4) — showing that (1) MetaMuse generates better cache replacement and bin packing solutions, outperforming LLM-based baselines (by up to 9.89% fewer cache misses, and up to 21.06% less bin usage) and human heuristics (by up to 35.76% fewer cache misses, and up to 30.93% less bin usage); (2) MetaMuse has up to 1.78 more distinct cache replacement solutions and 1.80 more distinct bin packing solutions, than LLM-based baselines; (3) MetaMuse has a low per-solution cost, up to 4.93 cents with GPT-4o.
2 Background and Motivation
2.1 System Algorithm Design
System algorithms define how computing systems behave. They are typically designed to optimize some performance objectives (e.g., cache hit ratio), for some scenarios or workloads (e.g., web servers). The difficulty of designing such algorithms arises from the nature of their solution space. It is discontinuous, where even a small change in algorithm design (e.g., the use of data structures) can lead to sharp and non-linear changes in performance. In addition, the discontinuity does not provide sufficiently predictable patterns or smooth landscape, to guide the search. This is a significant departure from prior efforts on auto-tuning system config parameters (Alipourfard et al., 2017; Cortez et al., 2017; Liang et al., 2020), which can be formulated as numerical optimization in most cases.
To this end, our work explores the use of LLMs to design heuristic algorithms, i.e., the algorithm generation task. Due to the discontinuous solution space, we approach this task as a sampling process in this space. At each step, LLMs attempt to generate distinct solutions. Conceptually, solving such a discontinuous task requires certain “Eureka” ideas that constitute leaps in the progress towards the final solution (Bubeck et al., 2023). Here, we refer to these leaps as creative ideation.
2.2 Creative Ideation
The goal of creative ideation is to discover useful solutions to a user-given problem, through a process of generating diverse solutions over time. The first requirement is usefulness — a generated solution should be relevant to the problem. The second requirement is diversity — a generated solution should produce feedback (or an outcome) that is unseen in the current process (Boden, 1998). As we accumulate diverse and useful solutions, the process advances towards finding the optimal solution.
For the algorithm generation task, creative ideation can be formulated as the following iterative process. Each iteration takes in problem statements of the algorithm, along with the set of functions to be implemented. For example, most caching algorithms can be abstracted into insert and evict functions (Yang et al., 2020). We can also take in metadata and feedback of previous solutions, as optional inputs. At iteration , the output is an executable cache solution () with all functions coded. We can optionally provide feedback to iteration , by measuring hit ratio of previous solutions (, …, ) on user-given workload trace, in an environment such as simulators. After cache solutions, we select the best-performing one from (, …, ).
2.3 LLM Limitations in Creative Ideation
Our work is motivated by the observation, where LLMs’ creative ideation capability is fundamentally limited by the same mechanism that enables it — the next-token prediction. In cognitive terms, responses are influenced by what one has frequently or recently seen, i.e., availability bias. For LLMs, this bias arises from the data frequency distribution in the training dataset.
Impacts on algorithm generation.
Due to this bias, LLMs tend to generate certain designs. We illustrate with cache replacement, using GPT-4o, Llama3.3-70B and DeepSeek-V3.
Figure 1 shows our attempts with Repeated Sampling (Brown et al., 2024; Snell et al., 2024). At iteration , each LLM is prompted to create a new design , with the default temperature of 1. Then, is implemented by GPT-4o in Python, and benchmarked in the simulator to measure hit ratios on 30 traces. Figure 1 clusters 1,200 solutions into centroids of human heuristics, according to the similarity of hit ratios. And, LLMs tend to design caches behaving similarly to well-known heuristics: least-recently used (LRU), least-frequently used (LFU), and first-in-first-out (FIFO). Another observation is that making Repeated Sampling aware of previous solutions does not mitigate this bias.
Is adjusting LLM hyperparameters all you need?
One such parameter is temperature (Ackley et al., 1985), which internally regulates randomness in the generative process. Temperature smooths the probability distribution of next-token candidates, as computed by the softmax of their logits. At higher temperatures, high probabilities are decreased, and low probabilities are increased.
Unfortunately, high temperatures only mitigate availability bias, without practically addressing it. The reason is that increasing temperature smooths the probability distribution, which is a monotonic transformation retaining the relative ranking of output token candidates. Even at extreme temperatures, the distribution would approach flat and result in incoherent LLM outputs.
2.4 Related Work
The first category is automatic heuristic design (Liu et al., 2024b; Ye et al., 2024; Zheng et al., 2025; Romera-Paredes et al., 2024; Novikov et al., 2025; Sharma, 2025). These efforts mostly rely on LLMs to evolve an initial population of candidate designs, through mutations and crossovers. This dependence on the initial population inherently constrains the reachable solution space. In addition, availability bias can impact how the population is evolved. Our case study highlights a bias towards frequently tweaking the cache algorithm’s scoring function, rather than leaping to other design dimensions. MetaMuse addresses these limitations by making leaps in the solution space.
There are also efforts exploring human-LLM collaborations to produce diverse outputs (Liu et al., 2024c; Wan et al., 2024; Vaccaro et al., 2024). However, they require human efforts and can be subject to human bias. Instead, MetaMuse posits that LLMs themselves are capable of thinking “outside the box” (Mehrotra et al., 2024), and proposes to combine three self-reflection principles to achieve this capability (§3).
3 MetaMuse
Towards effective creative ideation, we address availability bias through a principled framework of self-reflection, MetaMuse. The key is to enable LLMs to think “outside the box,” i.e., exploiting knowledge that otherwise seems probabilistically irrelevant to solving the user problem. Figure 2 shows the overview of an iteration, which outputs one executable solution. The first step is to evaluate the diversity of generated solutions so far (§3.1). This evaluation is based on the embedding computed from each solution’s feedback (e.g., cache hit ratios in the figure). The second step is to steer ideation, by selecting a set of stimuli (e.g., the “flower” keyword in the figure) based on the diversity evaluation. §3.2 describes two stimuli selection strategies in the current implementation: RSDict and RSDict-SF. Finally, MetaMuse develops stimuli into executable solutions (§3.3), through four waypoints of reasoning: property extraction, problem mapping, solution formulation, and code generation.
3.1 Evaluating Diversity
MetaMuse grounds diversity evaluation in feedback space, rather than idea space. In our case study, the former refers to performance measurements such as cache hit ratio, and the latter refers to the semantic embedding of algorithm descriptions or code implementations. Therefore, MetaMuse represents each solution with a feedback embedding, which is a vector of feedback measurements on different workload traces.
This principle is motivated by the observation that semantic differences do not necessarily imply distinct solutions. Considering a case where GPT-4o generated functionally equivalent LFU-based solutions twice in different descriptions: “The cache evicts the object that is least frequently accessed,” and “The cache evicts the object with the lowest priority. Upon a cache hit, the object’s priority is incremented.” Code semantics exhibit similar shortcomings, where a solution implemented with a linked list or a priority queue can have the same hit ratio.
Furthermore, feedback embedding has two desirable properties. First, since each dimension is a quantitative metric conveying magnitude, the Euclidean distance of two embeddings is a direct indication of how different corresponding solutions are. For the LFU case above, those two solutions would be considered equivalent because their distance is 0. Second, since feedback metrics typically pre-define a fixed range (e.g., 0–100% for cache hit ratio), the -dimensional embedding space is inherently structured and bounded. This allows us to compute the steering direction (§3.2).
3.2 Steering with External Stimuli
The objective of steering is to guide LLMs to generate solutions, which targets regions of the feedback embedding space. Rather than relying on LLMs’ internal randomness, MetaMuse uses external stimuli as the starting point for ideation. Stimuli should be unbiased to the problem, forcing LLMs to associate with knowledge that seems probabilistically irrelevant. One domain-agnostic instantiation of stimuli is keywords from an English dictionary.
Realizing this principle of steering has practical challenges. The first is the source of stimuli. In the case of keywords, a large dictionary offers a broader keyword coverage, but it also contains many keywords of little value. An example is technical jargon, e.g., “prolamin” in biology. In our case study, if LLMs are forced to associate these keywords, they would frequently use them as variable names, rather than inspiration to ideate new solutions. The second challenge is selecting stimuli at each iteration. §3.2.1 next presents selection strategies in our current implementation.
3.2.1 Stimuli Selection Strategies
Our current implementation of MetaMuse includes two strategies: RSDict and RSDict-SF.
RSDict.
RSDict is a stateless strategy that always randomly selects keywords from the dictionary. Since RSDict does not rely on evaluation results of previous solutions, it can steer scenarios where evaluation is infeasible or costly (e.g., human judges).
RSDict-SF.
Unlike RSDict, RSDict-SF considers the feedback embeddings of previous solutions, to compute the steering direction at each iteration. Steering direction is specified by a feedback embedding, which serves as the objective that RSDict-SF should aim to achieve in subsequent solutions. In our case study, steering direction enables both exploration (for diversity) and exploitation (for usefulness). For the former, the steering direction should be the farthest point from all previous solutions, i.e., the feedback embedding that has the greatest Euclidean distance from all other embeddings. For the latter, we set all dimensions of the target embedding vector to be a high value (e.g., 100% cache hit ratio).
Given an embedding of steering direction, RSDict-SF computes the set of stimuli that would likely develop into solutions close to it. Our current implementation formulates this step as a prediction problem, and solves it with Gaussian Process Regression (GPR) models. Given the sum of semantic embeddings of stimuli, we use GPR models (, …, ), to predict each of the dimensions in embedding. Having (, …, ) allows RSDict-SF to predict the expected feedback embedding of any set of stimuli. GPR models are fitted with all previous solutions at each iteration. We use the dot-product kernel for GPR models. Dot-product kernel can capture the pairwise semantic similarity between all observations across two solutions and is invariant to their ordering. For two solutions and , their similarity captured by is
where denotes the observations derived from ’s stimuli, and similar for .
We highlight two implementation details. First, predicting for all possible sets of stimuli is not feasible in practice, and a dictionary with 3,000 common English words would already have 3,000s sets. To this end, MetaMuse exercises the power-of-two random choices (Mitzenmacher et al., 2001). It randomizes two sets of stimuli from the dictionary, and predicts their expected feedback embeddings. Then, MetaMuse selects the one closest to the target embedding. Second, RSDict-SF bootstraps GPR models, with solutions generated by RSDict.
3.3 Developing Stimuli into Executable Solutions
Third, developing executable solutions from stimuli involves structured checkpoint-based steps, or waypoint reasoning. Waypoints are intermediate questions serving as checkpoints, which LLMs need to answer sequentially along the reasoning path. We observe that waypoints prevent LLMs from superficially developing solutions, e.g., simply turning stimuli into variable names in code. This addresses the problem with simply using free-form chain-of-thought (Wei et al., 2023; Mehrotra et al., 2024).
Our current implementation specifies the following waypoints. The first waypoint focuses on property extraction, where LLMs are prompted to reason about concepts and properties of the given stimuli. Considering the keyword “flower”, GPT-4o outputs “a ring of petals” as one property. The second waypoint focuses on problem mapping, and LLMs reason about problem-related observations that can be derived from these concepts and properties. In the previous example, GPT-4o outputs an observation that a cache can use a circular buffer. The third waypoint is solution formulation, which prompts LLMs to generate the complete description of a new solution. The observation derived from “flower” would be combined with other observations, for a complete solution. The final waypoint is code generation, and LLMs turn the solution description into executable Python code.
4 Empirical Results
Baselines.
We have 21 extensive baselines. LLM-based heuristics design baselines include MCTS-AHD (Zheng et al., 2025), ReEvo (Ye et al., 2024), OpenEvolve (Sharma, 2025), PlanSearch (Wang et al., 2025), and Repeated Sampling (Brown et al., 2024). We also include human-crafted SOTA heuristics for cache replacement problem (LRU, LFU, FIFO, Sieve (Zhang et al., 2024), S3FIFO (Yang et al., 2023), TinyLFU (Einziger et al., 2017), SLRU (Huang et al., 2013), Clock (Corbato, 1969), and ARC (Megiddo et al., 2003)), and online bin packing problem (Next Fit, Worst Fit, Almost Worst Fit (AWF) (Johnson, 1973), First Fit (Dósa & Sgall, 2013), Best Fit (Dósa & Sgall, 2014), Harmonic-k (Lee & Lee, 1985), and Refined First Fit (RFF) (Yao, 1980)). For Harmonic-k, we set =4 to align its number of bin categories with RFF.
MetaMuse.
We take a dictionary of common English words and remove stop-words to get 2,899 keywords. Each solution is ideated from =4 keywords. The feedback embedding consists of performance measurements on =30 traces. For cache replacement, these traces are generated by libCacheSim (Yang et al., 2020), from different Zipfian distributions. For bin packing, traces are from various Weibull distributions. The number of RSDict-SF warmup solutions is =100, roughly one-third of the total solutions in one experiment. Finally, prompts are included in §A.
Experiment setup.
We focus on two high-impact problems at a global cloud provider: cache replacement and bin packing. In each experiment, all methods aim to ideate and generate 350 executable solutions. To the best of our ability, we configure PlanSearch to output at least 15, 11, and 12 observations, while ideating with GPT-4o, Llama3.3-70B, and DeepSeek-V3, respectively. Many baselines such as MCTS-AHD can tune parameters in their solutions. §5 discusses safeguards against unsafe solutions, and environments are instrumented to catch errors, e.g., long-running execution. Unsafe solutions are re-implemented. Coding prompts are in §A.2.
To evaluate cache replacement solutions, we use 96 real-world workload traces from four data access scenarios: RetrievalAttention (Liu et al., 2024a) (24 “ra-fwe” and 24 “ra-multikey” traces), Tencent block storage (Zhang et al., 2020; 2018) (24 “tencent-storage” traces), and Alibaba cloud storage (Alibaba, 2020; Li et al., 2020; Wang et al., 2022) (24 “alibaba-storage” traces). The cache capacity is set to 10% of the number of distinct objects in each trace.
To evaluate online bin packing solutions, we use 288 workload traces: BPPLIB library (Delorme et al., 2018) (72 “Falkenauer-U” and 72 “Scholl-1” traces), Weibull distribution with parameter (shape=3, scale=45) (Castineiras et al., 2012) (72 “Weibull”), and Gaussian distribution with parameter (mean=0.3662, std=0.1416) (Yan et al., 2022) (72 “Gaussian”).
4.1 MetaMuse Generates High-performing Solutions
Cache replacement.
Figure 3 compares top solutions, as selected by the average cache miss ratio over all 96 workload traces. Box plots show their miss ratio reduction, with respect to FIFO heuristics.
We first delve into results from GPT-4o. At the 90th-percentile trace, MetaMuse achieves 5.17%–9.89% lower miss ratio than LLM-based baselines, and 1.75%–13.03% lower than human heuristics. At the 75th-percentile trace, MetaMuse achieves 3.62%–6.39% lower miss ratio than LLM-based baselines, and 6.62%–35.76% lower than human heuristics. We also see the improvement with Llama3.3-70B — at the 90th-percentile trace, MetaMuse achieves 2.41%–7.20% lower miss ratio than LLM-based baselines, and 0.94%–10.34% lower than human heuristics; at the 75th-percentile trace, MetaMuse achieves 3.67%–5.17% lower miss ratio than LLM-based baselines, and 5.48%–34.62% lower than human heuristics. On DeepSeek-V3, at the 90th-percentile trace, MetaMuse achieves up to 6.05% lower miss ratio than LLM-based baselines, and up to 9.14% lower than human heuristics.
Online bin packing.
Figure 4 compares top solutions, as selected by the average bin usage over all 96 workload traces. Box plots show their bin usage reduction, with respect to First Fit heuristics.
We first delve into results from GPT-4o. At the 90th-percentile trace, MetaMuse achieves 9.25%–9.42% lower bin usage than LLM-based baselines, and 9.25%–20.59% lower than human heuristics. At the 75th-percentile trace, MetaMuse achieves 6.94%–7.75% lower bin usage than LLM-based baselines, and 7.56%–18.36% lower than human heuristics. On Llama3-7B, at the 90th-percentile trace, MetaMuse achieves up to 0.19% less bin usage than LLM-based baselines, and up to 12.50% lower bin usage than human heuristics. On DeepSeek-V3, at the 90th-percentile trace, MetaMuse achieves up to 21.06% less bin usage than LLM-based baselines, and up to 30.93% lower bin usage than human heuristics.
4.2 MetaMuse Generates the Most Diverse Set of Solutions
Cache replacement.
The discovery of useful solutions depends on having diverse solutions. We evaluate diversity by the number of solutions with a distinct feedback embedding. Across all LLMs, MetaMuse consistently achieves higher diversity over LLM-based baselines. From analyzing 350 solutions generated by each method on GPT-4o, MetaMuse has 1.47 more distinct solutions on average. On DeepSeek-V3, MetaMuse has 1.57 more distinct solutions on average. On Llama3.3-70B, results show that MetaMuse can have 1.78 more distinct solutions on average.
Furthermore, we observe that having a higher diversity translates into lower availability bias. We recall that Repeated Sampling tends to generate solutions that tightly cluster to well-known human heuristics such as LRU, LFU, and FIFO, due to availability bias (§2.3). Using the nine human heuristics as centroids, empirical results show that MetaMuse reduces the cluster density (computed by ), for nearly all centroids. In other words, MetaMuse solutions are less tightly clustered than Repeated Sampling. For example, MetaMuse reduces the LRU cluster density by 35.25% (i.e., 6.15 down to 3.98). We see cluster density reduction even on the less common heuristics, e.g., the ARC cluster density reduces by 77.45% (i.e., 1.02 down to 0.23).
Online bin packing.
We evaluate diversity by the number of solutions with a distinct feedback embedding. Across all LLMs, MetaMuse consistently achieves a higher diversity over LLM-based baselines. From analyzing 350 solutions generated by each method on GPT-4o, MetaMuse has 1.44 more distinct solutions on average. On DeepSeek-V3, it has 1.80 more distinct solutions on average. On Llama3-70B, it has 1.31 more distinct solutions on average.
4.3 MetaMuse Has a Low Per-solution Cost
We see having a low per-solution cost, as a factor driving the adoption of LLMs for creative ideation. We illustrate with the case of generating cache replacement solutions. The total cost for MetaMuse to generate one full solution is up to 4.93 cents with GPT-4o, up to 3.80 cents with DeepSeek-V3, and up to 3.85 cents with Llama3.3-70B. Looking at the cost breakdown, it costs 3.72 cents for GPT4-o to output one implementation in Python. We note that if an implementation is considered unsafe by our safeguards (§5), we re-implement it.
4.4 Surprising Designs from MetaMuse
In addition to quantitative evaluations, we highlight surprising designs not immediately obvious to engineers. We consider a top-performing cache replacement solution: MetaMuse-533 (§B.1).
One is the NSE counter tracking the number of times that a cached object has witnessed eviction events. It contributes to the score calculation, along with recent access time and frequency counter. Due to NSE, MetaMuse-533 behaves to favor objects that have remained longer in the cache (up to some thresholds). Interestingly, NSE seems to counter the common belief of favoring newly admitted objects, in order to learn their access patterns and prevent thrashing (Denning, 1968).
Another is the use of saturating counters, which increment until some pre-defined threshold. The typical usage is to reduce storage overhead by capping the maximum value of a counter (Yang et al., 2023; Corbato, 1969). Instead, MetaMuse-533 seems to use saturating counters to accumulate meaningful usage history, but not so much as to mislead eviction decisions. For example, a saturating frequency counter allows MetaMuse-533 to identify bursty objects, while preventing their counts from growing too large during the burst.
4.5 Ablation Study
We test key components in MetaMuse, with the case study of cache replacement algorithms.
Stimuli selection strategies.
We test RSDict and RSDict-SF by running experiments on GPT-4o. First, the majority of RSDict and RSDict-SF solutions outperform Repeated Sampling solutions, and this demonstrates the value of having stimuli for steering ideation. Comparing solutions for the 90th-percentile workload trace, 67.20% of RSDict solutions and 75.60% of RSDict-SF solutions have a higher cache hit ratio than Repeated Sampling. In addition, these hit ratios can be up to 13.13% and 27.07% higher, respectively. Similarly, for the 75th-percentile workload trace, we see 66.80% of RSDict solutions and 72.40% of RSDict-SF solutions outperform Repeated Sampling. In addition, these hit ratios can be up to 15.43% and 21.93% higher, respectively.
Second, compared to RSDict, RSDict-SF has a greater number of solutions that can outperform Repeated Sampling. Empirical results show 12.50% and 8.38% more solutions, with respect to the 90th percentile and 75th percentile of workload traces, respectively. Furthermore, looking at solution diversity, RSDict-SF results in 13.17% more distinct solutions than RSDict.
Waypoint reasoning.
Figure 5 shows that waypoint reasoning enables MetaMuse to generate more diverse solutions. We fix the coding agents to GPT-4o. At each iteration, we instrument MetaMuse to generate solutions with and without waypoint reasoning. When ideating with GPT-4o (5(a)), waypoint reasoning improves the number of distinct solutions from 149 to 175 for RSDict, and from 152 to 197 for RSDict-SF. When ideating with DeepSeek-V3 (5(c)), waypoint reasoning improves the number of distinct solutions from 113 to 144 for RSDict, and from 119 to 140 for RSDict-SF. When ideating with Llama3.3-70B (5(b)), waypoint reasoning improves the number of distinct solutions from 124 to 148 for RSDict, and from 129 to 154 for RSDict-SF.
5 Discussion
Safeguards against unsafe solutions.
In the case of system algorithms, three main errors result in unsafe implementations: (1) long-running execution (e.g., more than 5 seconds), (2) excessive memory usage, and (3) illegal behaviors, such as falsely claiming a cache hit when the requested object is not present. To this end, our environments are configured to monitor the process execution time and memory usage. Furthermore, we constrain the coding scope for LLMs by explicitly specifying the functions (e.g., insert and evict for caching algorithms). Finally, with Python scripts as outputs, we can leverage testing and verification tools in the Python ecosystem.
Limitations and future discussion.
First, our implementation explores an instantiation of external stimuli (§3.2), which is specifically chosen for being general and problem-agnostic. However, the MetaMuse framework is extensible to take in other instantiations, and future work will explore possibilities such as long-term task-related memories. Second, our work currently focuses on fully automated ideation, and an interesting direction is to explore whether subjective human guidance (e.g., human expertise) could further improve the solution usefulness. Third, we evaluate on real-world workload traces (§4), and deploying LLM-generated algorithms to a business-critical production environment remains an ongoing collaboration with engineers at a global cloud provider. We look forward to leverage lessons learned for deeper practical impacts.
6 Conclusion
To practically drive algorithm generation, MetaMuse is a framework of creative ideation that overcomes LLMs’ availability bias. MetaMuse combines three self-reflection principles, guiding the stages of creative ideation. Evaluations show that it can generate high-performing solutions for two high-impact problems at a global cloud provider: cache replacement and bin packing algorithms. Furthermore, we observe surprising design considerations not immediately obvious to human engineers.
Acknowledgments
We thank Encheng Xie (University of Illinois Urbana-Champaign) for his valuable discussions and contributions to the experiments.
Appendix A Prompts
A.1 Ideation Prompts
Repeated Sampling uses the following prompt for solution formulation, generating its natural language description:
We use PlanSearch’s original prompts (Wang et al., 2025) to generate observations. These observations are then inserted into the placeholder [[hints]] of the following prompt for solution formulation:
MetaMuse uses the following prompt for property extraction and problem mapping. They are the first two waypoints (c.f. §3.3).
For solution formulation, MetaMuse uses the same ideation prompt as PlanSearch. The placeholder [[hints]] will be replaced with the observations derived from the first two waypoints.
A.2 Coding Prompt
All methods use the following prompt to implement a solution in Python. The placeholders, [[metadata]], [[evict]], [[update_after_hit]], [[update_after_insert]], and [[update_after_evict]], will be filled in with the corresponding fields in the JSON-formatted output of the ideation prompts. [[design]] will be filled with a concatenation of these fields.